OpenStudy (bossimbacon):

What is the slope of the line that passes through the points (1, 3) and (5, -2)? (Points : 4) Option A: 5/4 Option B: -5/4 Option C: 4/5 Option D: -4/5

OpenStudy (alexandervonhumboldt2):

use slope formula: \[\frac{ y_1-y_2 }{ x_1-x_2 }=\frac{ 3-(-2) }{ 1-5 }=?\]

OpenStudy (bossimbacon):

so its b? @AlexandervonHumboldt2

OpenStudy (alexandervonhumboldt2):

i got B too